# ReLaX-VQA: Residual Fragment and Layer Stack Extraction for Enhancing Video Quality Assessment

### v1

@@ Abstract
With the rapid growth of User-Generated Content (UGC) exchanged between users and sharing platforms, the need for video quality assessment in the wild has emerged. UGC is mostly acquired using consumer devices and undergoes multiple rounds of compression or transcoding before reaching the end user. Therefore, traditional quality metrics that require the original content as a reference cannot be used. In this paper, we propose ReLaX-VQA, a novel No-Reference Video Quality Assessment (NR-VQA) model that aims to address the challenges of evaluating the diversity of video content and the assessment of its quality without reference videos. ReLaX-VQA uses fragments of residual frames and optical flow, along with different expressions of spatial features of the sampled frames, to enhance motion and spatial perception. Furthermore, the model enhances abstraction by employing layer-stacking techniques in deep neural network features (from Residual Networks and Vision Transformers). Extensive testing on four UGC datasets confirms that ReLaX-VQA outperforms existing NR-VQA methods with an average SRCC value of 0.8658 and PLCC value of 0.8872. We will open source the code and trained models to facilitate further research and applications of NR-VQA: https://github.com/xinyiW915/ReLaX-VQA.
1 Introduction
@@
Video Quality Assessment (VQA) is critical for optimising user experience in video streaming. However, due to the explosion of User-Generated Content (UGC) on video sharing platforms such as YouTube and TikTok sandvine2024 , VQA has become increasingly challenging. In UGC use cases, videos are typically first encoded on the user’s device (smartphones, consumer cameras, etc) while being captured, and then transcoded by the platform for distribution and storage. As a result of repeated lossy compression, video quality is degraded by the introduction of noticeable artefacts such as blocking, ringing, and blurring Adsumilli2019 ; wang2019youtube .
@@
Traditional VQA methods that rely on comparing the distorted video with a reference video, using measures such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) wang2004image , cannot be reliably used for UGC, as the uploaded video to the sharing platform is not the original. Similarly, learning-based full reference methods, such as Video Multi-Method Assessment Fusion (VMAF)li2016toward , cannot capture the quality of transcoded content because it has been trained on pristine original references and only once-compressed content. Therefore, developing No-Reference Video Quality Assessment (NR-VQA) models for evaluating these types of no-reference video content is imperative.
@@
Over the past few years, Deep Neural Networks (DNNs) have been used to devise advanced NR-VQA models. The robust representation learning capability of DNNs has positioned them as the prime technique for tackling visual tasksalzubaidi2021review . Such models usually employ 2D-CNN, 3D-CNN, and Transformersliu2018end ; li2019quality ; tu2021rapique ; ying2021patch ; sun2022deep ; wu2023discovqa ; zhao2023zoom and are designed to analyse pixel variations due to compression and correlations to subjective quality scores. Existing models are primarily frame-based and struggle to capture temporal impairments effectively. Furthermore, the resizing of frames causes further degradation, which has driven the development of VQA models focused on fragmentswu2022fast ; zhao2023zoom . Despite their efficiency, DNN VQA models are resource-thirsty (high computational complexity) and data-hungry (requiring large volumes of data for training) sze2017efficient . Thus, the unavailability of large-scale datasets with subjective scores to train these networks from scratch limits their performance. Therefore, many existing VQA models rely on networks that are pre-trained on large image classification datasets to extract frame-based featuresli2019quality ; tu2021rapique ; you2021long ; madhusudana2023conviqt . Although results look promising, this practice can result in a distributional shift between the original training task and the actual VQA application task. Moreover, despite the importance of frame-level spatial information, temporal information (motion or inter-frame variations) needs to be accounted to evaluate the video quality accurately.
@@
In response to these challenges, we introduce ReLaX-VQA, a feature extractor that utilises two main techniques: spatio-temporal fragmentation (residual and optical flow) and layer stacking. In consideration of the residual between successive frames, inspiration is drawn from the structure of current block-based video encoders (e.g., H.264 r:h264 or H.265 r:HEVC ), where the frame encoding follows a hierarchical structure within a group of frames, with more bits allocated for the encoding of keyframes (I frames) and fewer for the successive predicted frames (P frames). Thus, the quality of the decoded frames varies per frame and is typically higher for I frames. Therefore, when subtracting two successive frames, key and a predicted, the residual is an expression of both the temporal information in moving regions as well as spatial artefacts around areas with compression artefacts. This residual information can be exploited to correlate it with the perceptual variations in the Human Visual System (HVS). As this residual is sparse, we propose the extraction of residual fragments. The selection of the fragments is based on the ranking of residuals (absolute difference) on the basis that human attention is drawn to areas with visible distortions. Fragmenting is also applied to temporal information, particularly to optical flow, which can reveal regions in the frame undergoing the most dynamic changes. This method can effectively utilise the motion information to enhance the perception of the model. We also selected patches from locations of prominent dynamic changes in the sample frame. The second technique relies on a layer-stacking DNN feature extraction. Unlike pooling, this enhances the representation of deep features and increases their relevance to human perception. Through visualisation and perceptual correlation, we identified that different network layers capture information at different levels, from edges to abstract features. This hierarchical feature extraction mechanism allows our model to mimic human visual processing more closely, thereby achieving more precise visual perception in complex scenarios.
@@
Finally, to further improve the prediction performance of the extracted features, we constructed a simple and efficient Multi-Layer Perceptron (MLP) regression head for training and testing. By combining the characteristics of the Residual Network (ResNet) he2016deep focusing on local information, and the advantages of the Vision Transformer (ViT) vaswani2017attention which concentrates on global features, our method improves the prediction accuracy. The contributions of this work are summarized below:
  1. 1.
    We propose ReLaX-VQA, a novel NR-VQA model that focuses on extracting fragmented information from frame differencing and optical flow fields between consecutive frames to enhance spatio-temporal perception.
  2. 2.
    We boost the feature abstraction capabilities by layer-stacked feature extraction. The layer stacking technique has proven effective in information fusion.
  3. 3.
    We provide an improved design of MLP regression that fuses local and global features for higher accuracy.
@@
The proposed model, ReLaX-VQA, has undergone extensive testing on several public large-scale datasets. The results, assessed by linear and rank correlation coefficients with mean opinion scores, Spearman (SRCC), Kendal (KRCC), and Pearson (PLCC), as well as Root Mean Square Error (RMSE), indicate higher accuracy, outperforming existing NR-VQA methods. Our model is straightforward to implement and computationally efficient. We will make the source code and pre-trained models available for public research and use.
2 Related work
Table 1: Summary of VQA benchmark datasets
@@ 2 Related work
CVD2014nuutinen2016cvd2014 480p, 720p 10-25s In-capture In-lab 234
KoNViD-1khosu2017konstanz 540p 8s In-the-wild Crowdsourced 1,200
LIVE-VQCsinno2018large 240p-1080p 10s In-the-wild Crowdsourced 585
YouTube-UGCwang2019youtube 360p-4k 20s In-the-wild Crowdsourced 1,149*
LSVQying2021patch 99p-4k 5-12s In-the-wild Crowdsourced 38,793**
2.1 UGC video datasets
@@
To develop more realistic and accurate VQA models for UGC content, numerous datasets have been created. Table 1 reports commonly used subjective video quality datasets used in recent studies. CVD2014 nuutinen2016cvd2014 and LIVE-VQC sinno2018large are datasets that simulate distorted videos from online video streaming through compression and transmission distortion, and artificial distortions (e.g., Gaussian blurring)dodge2016understanding . These datasets are small and contain only a limited number of unique source videos with restricted content diversity and distortion complexity. On the other hand, video quality databases collected “in the wild” specifically for UGC are constantly being updated. These collections include a large number of real distorted videos produced by amateurs and more accurately reflect the complexities of UGC videos. Examples are the LIVE-Qualcommghadiyaram2017capture , KoNViD-1khosu2017konstanz , YouTube-UGCwang2019youtube , and LIVE-VQCsinno2018large . Although these datasets are larger and contain more content variation, they are still insufficient to train deep learning models. The recent release of two large-scale VQA datasets: FlickrVid-150kgotz2019no (not publicly available) and LSVQying2021patch tried to address the lack of training data for deep learning-based VQA models. Notably, the publicly available LSVQ dataset, which contains 39,076 videos, demonstrates advantages in handling large-scale video data.
2.2 NR VQA models
@@
Numerous classic NR-VQA models have been developed based on traditional Natural Scene Statistics (NSS), such as NIQEmittal2012making , PIQEmoorthy2011blind , and BRISQUEmittal2011blind . Inspired by frequency domain characteristics, V-BLIINDS6705673 and V-CORNIA7025098 used Discrete Cosine Transform (DCT) statistics and learned codebooks to mark image integrity and block features respectively. These methods perform better for specific types of distortions (e.g., blurriness) and have been designed to capture a specific range of video quality. Other models employ a large number of handcrafted features and feature selection strategies to evaluate video quality. Successful examples are TLVQMkorhonen2019two and VIDEVALtu2021ugc . TLVQM combines high spatial complexity and low temporal complexity features, while VIDEVAL integrates a wide range of handcrafted features to simulate real distortions. These models train shallow regression models with statistical features and perform well on standard video (not UGC) content, but are limited in generalising over UGC content.
@@
More recently, VQA research has shifted towards DNNs, such as CNNs and Transformers for constructing feature extractors. Notably, RAPIQUE tu2021rapique , combines NSS with semantic-aware CNN features extracted from downsampled frames, resulting in reduced computational runtime. Another high-performing model is VSFAli2019quality and its enhanced version, MDTVSFAli2021unified (using a hybrid dataset training strategy), that extracts frame features using the pre-trained image classification model ResNet-50he2016deep and apply gated recursive units to process temporal features. The latest developments in NR-VQA focus on addressing multi-resolution and complex spatio-temporal distortions in UGC. Using patches and fragments as in ying2021patch ; wu2022fast ; zhao2023zoom and/or novel data sampling mechanisms as in liu2024scaling ; ke2023mret helps preserve the global and local details of high-resolution videos, capturing complex spatio-temporal features and improving processing efficiency. Other NR-VQA models perform end-to-end feature training or multi-scale feature fusion, utilising attention mechanisms and simulating HVS features to enhance performancewu2022fast ; lao2022attentions ; zhang2022hvs ; sun2022deep . An important aspect of VQAs is the ranking ability and therefore Chen et al Feng_2024_WACV focus on improving the ranking performance of their model by using VMAF-based quality ranking information but train their model on standard video datasets with access to uncompressed content. Li et al.li2022blindly address domain bias by transferring knowledge from image quality and motion recognition databases to improve benchmarks. UVQwang2021rich further augments the video quality metrics by incorporating semantic information.
3 Proposed method
@@
An overview of the proposed ReLaX-VQA model framework is provided in Fig. 1 and consists, from bottom to top, of the spatio-temporal fragment sampling module (Section 3.1), DNN Layer stack module (Section 3.2), and quality regression module (Section 3.3). Firstly, we extracted features from successive sampled video frames using the spatio-temporal fragment sampling module from frame differencing (residuals) and optical flow. Then, we stacked the extracted features using different layers of the deep neural network, fusing these multi-layered features to capture the quality-aware features. Finally, a simple MLP quality regression module maps the quality-aware features to video quality scores. Furthermore, we introduced a hybrid Mean Absolute Error (MAE) and Rank loss function to optimise the quality prediction model. Through this approach, ReLaX-VQA can effectively utilise the residual information after dynamic changes between video frames. The local and global information that affects the video quality is captured through feature layer stacking, enhancing the model’s abstraction capability and its consistency with human perception.
Figure 1: Overview of the proposed ReLaX-VQA framework that demonstrates the different processing of the input frames for the feature extraction and regression to infer the quality score. The architectures of ResNet-50 Stack (I) and ResNet-50 Pool (II) are provided in Fig. 2 in the Appendix.
3.1 Spatio-temporal fragment sampling module

### v3

# Xinyi Wang    Angeliki Katsenou    David Bull

@@ Abstract
With the rapid growth of User-Generated Content (UGC) exchanged between users and sharing platforms, the need for video quality assessment in the wild is increasingly evident. UGC is typically acquired using consumer devices and undergoes multiple rounds of compression (transcoding) before reaching the end user. Therefore, traditional quality metrics that employ the original content as a reference are not suitable. In this paper, we propose ReLaX-VQA, a novel No-Reference Video Quality Assessment (NR-VQA) model that aims to address the challenges of evaluating the quality of diverse video content without reference to the original uncompressed videos. ReLaX-VQA uses frame differences to select spatio-temporal fragments intelligently together with different expressions of spatial features associated with the sampled frames. These are then used to better capture spatial and temporal variabilities in the quality of neighbouring frames. Furthermore, the model enhances abstraction by employing layer-stacking techniques in deep neural network features from Residual Networks and Vision Transformers. Extensive testing across four UGC datasets demonstrates that ReLaX-VQA consistently outperforms existing NR-VQA methods, achieving an average SRCC of 0.8658 and PLCC of 0.8873. Open-source code and trained models that will facilitate further research and applications of NR-VQA can be found at: https://github.com/xinyiW915/ReLaX-VQA.
Index Terms:
No-Reference Video Quality Assessment, User-generated Content, Deep Features, Residual, Optical Flow, Video Compression
I Introduction
@@
Video Quality Assessment (VQA) is a critical component when optimising user experience in video streaming applications. However, due to the vast volumes of User-Generated Content (UGC) on video-sharing platforms such as YouTube and TikTok [1], VQA is challenging. In the UGC use case, videos are typically first encoded on a user device (smartphone, consumer camera, etc.) when captured and then are transcoded by the streaming platform for storage and distribution. Video parameters such as spatial resolution are often also modified to meet different user device or bandwidth constraints. As a result, UGC video is more susceptible to degradation (via noticeable artefacts such as blocking, ringing, and blurring) than professional content [2, 3].
@@
Traditional VQA methods that compare a distorted video with its pristine reference (e.g.) Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [4]) cannot be reliably used for UGC assessment, because the video uploaded to the sharing platform is not a pristine version. Similarly, learning-based full reference methods, such as Video Multi-Method Assessment Fusion (VMAF) [5], cannot capture the quality of transcoded content because they have been designed and trained on pristine original references compared to content that has only been compressed once. As a consequence, there is an urgent need to develop robust No-Reference Video Quality Assessment (NR-VQA) models.
@@
In recent years, Deep Neural Networks (DNNs) have been used to devise advanced NR-VQA models. The robust representation learning capability of DNNs has positioned them as the approach of choice for tackling a wide range of visual tasks [6]. Such models usually employ 2D-CNNs [7, 8, 9], 3D-CNNs [10, 9, 11], and Transformers [12, 13] and are designed to analyse pixel variations due to compression based on correlations with subjective quality scores. Existing models are primarily frame-based and struggle to capture temporal impairments effectively. Furthermore, the resizing of frames can result in further quality degradation, which has driven the development of VQA models that do not resize the frames but rather select frame fragments [14, 13]. Despite their efficiency, DNN VQA models are associated with high computational complexity and are data-hungry (requiring large volumes of data for training) [15]. Thus, the unavailability of large-scale datasets with subjective scores to train these networks from scratch limits their performance. Therefore, many existing VQA models rely on networks that are pre-trained on large image classification datasets to extract frame-based features [7, 8, 16, 17]. Although the results look promising, this practice can result in a distributional shift between the original training task and the actual VQA application task. Moreover, despite the importance of frame-level spatial information, temporal information (motion or inter-frame variations) needs to be accounted for to evaluate the video quality accurately.
@@
In response to these challenges, we introduce ReLaX-VQA, a feature extractor that utilises two main techniques: (i) selective spatio-temporal fragmentation based on frame differences (or residuals) and optical flow and (ii) feature layer stacking. For the frame differences, inspiration is drawn from the structure of current hybrid block-based video encoders (e.g., H.264 [18] or H.265 [19]) where, for most coding modes frame encoding follows a hierarchical structure within a group of frames, with more bits allocated for the encoding of keyframes (I frames) and fewer for the successive predicted frames (P frames). Thus, the quality of the decoded frames varies per frame and is typically higher for I frames. Therefore, when subtracting a predicted from a key frame, the residual is an expression of both the temporal information in moving regions as well as spatial artefacts around stationary areas with compression artefacts. This residual information can be exploited to correlate it with the expected regions of visual attention that reflect perceptual variations in the Human Visual System (HVS). As this residual is sparse, we propose the selective extraction of fragments. The residual frames are split with a grid, and the selection is based on the ranking of residual fragments (using the absolute difference). This was designed on the basis that human attention is drawn to areas with visible distortions. Additionally, fragmenting is also applied to temporal information, and particularly to the optical flow (OF) computed between the key and predicted frames. The OF can reveal regions within the frame with the most dynamic changes. This method can effectively utilise the spatio-temporal information to enhance the perception of the model. Based on the OF fragment selection, we also selected fragments from the corresponding locations in the sample video frame.
The second technique relies on layer-stacking DNN feature extraction. Unlike deploying pooling techniques at a layer level that produces abstract representations, stacking features enhance the representation of deep features and increase their relevance to multi-layered human perception. Through visualisation and perceptual correlation, we identified that different network layers capture information at different levels, from edges and textures to abstract features. This hierarchical feature extraction mechanism allows our model to mimic human visual processing more closely, thereby achieving more precise visual perception in complex scenarios.
@@
Finally, to further improve the prediction performance of the extracted features, we constructed a simple and efficient Multi-Layer Perceptron (MLP) regression head for training and testing. By combining the characteristics of the Residual Network (ResNet) [20], focusing on local information and the advantages of the Vision Transformer (ViT) [21], which concentrates on global features, our method improves the prediction accuracy.
The primary contributions of this work are summarized below:
  1. i.
    ReLaX-VQA Framework: We proposed ReLaX-VQA, a novel NR-VQA model that employs two different strategies for feature extraction: selective fragmentation and layer stacking. Extensive evaluation of our model demonstrates higher accuracy compared to other state-of-the-art metrics while maintaining a relatively low computational complexity (see Section IV).
  2. ii.
    Selective Feature Extraction: We focus on selectively extracting ranked fragments from frame differences of key and predicted frames, from OF fields, and from the key frame. This biases the extraction of features from areas that are more prone to transcoding or scaling errors, thus enhancing spatio-temporal perception. Although the idea of fragmentation for NR-VQA has been used before [14], our implementation significantly differs through a dynamic selection based on ranked spatio-temporal differences between frames (further explained in Section II) and delivers improved results.
  3. iii.
    Incorporating Layer Stacking: We boosted the feature abstraction capabilities by layer-stacked feature extraction from scaled (resized) frames using both ResNet and ViTs. The ResNets are better for a global feature representation, while ViTs have proven more effective for attention to salient regions. The layer stacking technique has proven effective in information fusion.
  4. iv.
    Improved MLP Regression: We have improved the design of MLP regression that fuses local and global features for higher accuracy. We designed different MLP regressors for large-scale and small datasets for prediction. Moreover, we used an SGD optimiser and applied Cosine Annealing learning rate decay and Stochastic Weight Averaging to optimize the training process.
@@
The proposed model, ReLaX-VQA, has been tested on several public large-scale datasets. The results, assessed by linear and rank correlation coefficients with mean opinion scores, Spearman (SRCC), Kendal (KRCC), and Pearson (PLCC), as well as Root Mean Square Error (RMSE), indicate higher accuracy on most datasets and consistently very high accuracy and correlation across datasets, outperforming existing NR-VQA methods. We have made the source code and pre-trained models available for public research and use.
The remainder of this paper is organized as follows. Section II summarises the related work in terms of existing datasets and NR-VQAs. Next, Section III describes the proposed framework. Then, Section IV details the experimental set-up, configurations, and results. Finally, Section V concludes this text with a summary of our findings and potential future directions.
Figure 1: Overview of the proposed ReLaX-VQA framework that demonstrates the different processing of the input frames for feature extraction and regression to infer the quality score. The visualisation of the intermediate representation of fragmented data is illustrated. More details on the architectures of ResNet-50 Stack (I) and ResNet-50 Pool (II) are provided in Fig. 2.
II Related work
TABLE I: Summary of VQA benchmark datasets
@@ II Related work
CVD2014 [22] 480p, 720p 10-25s In-capture In-lab 234
KoNViD-1k [23] 540p 8s In-the-wild Crowdsourced 1,200
LIVE-VQC [24] 240p-1080p 10s In-the-wild Crowdsourced 585
YouTube-UGC [3] 360p-4k 20s In-the-wild Crowdsourced 1,149*
LSVQ [9] 99p-4k 5-12s In-the-wild Crowdsourced 38,793**
II-A UGC video datasets
@@
To develop more realistic and accurate VQA models for UGC content, numerous datasets have been created. Table I reports commonly used subjective video quality datasets used in recent studies. CVD2014 [22] and LIVE-VQC [24] are datasets that simulate distorted videos from online video streaming through compression, transmission distortion, and artificial distortions (e.g., Gaussian blurring) [25]. These datasets are small and contain only a limited number of unique source videos with restricted content diversity and distortion complexity. On the other hand, video quality databases collected “in the wild” specifically for UGC, are constantly being updated. These collections include a large number of real distorted videos produced by amateurs and more accurately reflect the complexities of UGC videos. Examples are the LIVE-Qualcomm [26], KoNViD-1k [23], YouTube-UGC [3], and LIVE-VQC [24]. Although these datasets are larger and contain more content variation, they are still insufficient to train deep learning models. The recent release of two large-scale VQA datasets, FlickrVid-150k [27] (not publicly available) and LSVQ [9] tried to address the lack of training data for deep learning-based VQA models. Notably, the publicly available LSVQ dataset, which contains 39,076 videos, demonstrates advantages in handling large-scale video data.
II-B NR VQA models
@@
Numerous classic NR-VQA models have been developed based on traditional Natural Scene Statistics (NSS), such as NIQE [28], PIQE [29], and BRISQUE [30]. Inspired by frequency domain characteristics, V-BLIINDS [31] and V-CORNIA [32] used Discrete Cosine Transform (DCT) statistics and learned codebooks to mark image integrity and block features, respectively. These methods perform better for specific types of distortions (e.g., blurriness) and have been designed to capture a specific range of video quality. Other models employ a large number of handcrafted features and feature selection strategies to evaluate video quality. Successful examples are TLVQM [33] and VIDEVAL [34]. TLVQM combines high spatial complexity and low temporal complexity features, while VIDEVAL integrates a wide range of handcrafted features to simulate real distortions. These models train shallow regression models with statistical features and perform well on standard video (not UGC) content but are limited in generalising over UGC content.
@@
More recently, VQA research has shifted towards DNNs, employing CNNs and Transformers for constructing feature extractors. Notably, RAPIQUE [8] combines NSS with semantic-aware CNN features extracted from downsampled frames, resulting in reduced computational runtime. Other high-performing models include VSFA [7] and its enhanced version, MDTVSFA [35] (using a hybrid dataset training strategy). These extract frame features using a pre-trained image classification model ResNet-50 [20] and apply gated recursive units to process temporal features. The latest developments in NR-VQA focus on addressing multi-resolution and complex spatio-temporal distortions in UGC. Using patches and fragments as in [9, 14, 13] and/or novel data sampling mechanisms as in [36, 37] helps preserve the global and local details of high-resolution videos, capturing complex spatio-temporal features and improving processing efficiency. Other NR-VQA models perform end-to-end feature training or multi-scale feature fusion, utilising attention mechanisms and simulating HVS features to enhance performance [14, 38, 39, 11]. A powerful aspect of VQA metrics is the ability to rank quality levels. Chen et al [40] focus on improving the ranking performance of their model by using VMAF-based quality ranking information but train their model on standard video datasets with access to uncompressed content. Li et al. [41] address domain bias by transferring knowledge from image quality and motion recognition databases to improve benchmarks. UVQ [42] further augments the video quality metrics by incorporating semantic information.
II-C Advancing the State-of-the-Art
Two of the best-performing learning-based approaches mentioned above that were designed for the quality assessment of UGC content are VSFA [7] and Fast-VQA [14, 43]. These will comprise the benchmarks of the learning based methods for our experiments. VSFA [7] utilises content-aware features of pre-trained neural networks and gated recursive units to model temporal dependencies. However, the fixed feature extraction method may not fully capture the diverse quality variations of UGC content. In contrast, ReLaX-VQA embodies dynamic fragment selection based on frame differences to more effectively capture temporal variations. This enhances robustness and accuracy through deep learning feature extraction. ReLaX-VQA achieves higher linear and rank correlation, aligning more closely with human perception.
Wu et al. proposed Fast-VQA and DOVER models[14, 43, 44] that apply fragmentation using a uniform grid to randomly select fragments with the benefit of reduced computational costs. However, the uniform sampling strategy for fragmentation may omit perceptually important spatio-temporal changes in UGCs. Compared to Fast-VQA, DOVER provides the option of multimodality through an aesthetics perspective. ReLaX-VQA further advances this technique by employing a ranking of the spatio-temporal fragments based on pixel differences between key and predicted frames, using mean absolute differences to capture significant quality variations. As a result, ReLaX-VQA enhances both temporal and spatial perception, demonstrating higher accuracy and better alignment with human perception compared to existing NR-VQA methods. The results reported in Section IV-B confirm the effectiveness of our proposed method.
III Proposed method
@@
An overview of the proposed ReLaX-VQA model framework is provided in Fig. 1 and consists of three basic modules: the Spatio-Temporal Fragment Sampling module (Section III-A), the DNN Feature Extraction module (Section III-B), and Quality Regression module (Section III-C). Firstly, we extract features from successively sampled video frames using the spatio-temporal fragment sampling module: frames, frame differences (residuals), and optical flow. We also, resize the frames and extract DNN features from the entire frame. Then, we stack the extracted features from the resized and fragmented frame using different layers of the ResNet-50, thus fusing these multi-layered features to capture the quality-aware features. Here, we note that since the ViT is based on a self-attention mechanism, we do not use a layer-stacking framework for this, as also shown in Fig. 2. Finally, a simple MLP quality regression module maps the quality-aware features to video quality scores. Furthermore, we introduced a hybrid Mean Absolute Error (MAE) and Rank loss function to optimise the quality prediction model. Through this approach, ReLaX-VQA can effectively utilise the residual information after dynamic changes between video frames. The local and global information that affects the video quality is captured through feature layer stacking, enhancing the model’s abstraction capability and its consistency with human perception.
Figure 2: The bespoke architectures of ResNet-50 Stack (I), ResNet-50 Pool (II), and ViT Pool.
III-A Spatio-Temporal Fragment Sampling module